# Nuclear rainbow of the symmetric nucleus-nucleus system: Interchange of the nearside and farside scattering

### v1

@@ Abstract
Extensive elastic scattering data measured at energies around 10 to 20 MeV/nucleon for some light identical systems, like 12C+12C and 16O+16O , were shown to exhibit the nuclear rainbow pattern of a broad Airy oscillations of the elastic scattering cross section at medium and large angles. Because of the identity of the scattered projectile and recoiled target, the smooth rainbow pattern at angles around and beyond $\theta_{\rm c.m.}\approx 90^{\circ}$ is strongly deteriorated by the boson exchange in the 12C+12C and 16O+16O systems at low energies. The exchange symmetry of two identical nuclei implies the Mott interference of the direct and exchange scattering amplitudes, which destroys the nuclear rainbow pattern. The nuclear rainbow features in the elastic scattering of two identical nuclei have been discussed so far based on the nearside-farside (NF) decomposition of the scattering amplitude given by an optical model calculation neglecting the projectile-target exchange symmetry. Moreover, the NF decomposition method was developed in the seventies by Fuller for nonidentical dinuclear systems only, and the details of how the exchange symmetry of an identical system affects the evolution of nuclear rainbow remain unexplored. For this purpose, the Fuller method is generalized in the present work for the elastic scattering of two identical (spin-zero) nuclei, with the projectile-target exchange symmetry taken explicitly into account. The results obtained for elastic 12C+12C and 16O+16O scattering at low energies show that the exchange symmetry results in a symmetric interchange of the nearside and farside scattering patterns at angles passing through $\theta_{\rm c.m.}=90^{\circ}$, which requires a more subtle interpretation of nuclear rainbow. We found further that a similar NF interchange also occurs in a nonidentical nucleus-nucleus system with the core-core symmetry at low energies, where the elastic cross section at backward angles is due mainly to the elastic transfer of cluster or nucleon between two identical cores. This interesting effect is illustrated in the elastic 16O+12C scattering at low energies where the elastic $\alpha$ transfer between two 12C cores has been proven to enhance the elastic cross section at backward angles.
@@ I Introduction
To motivate the present study, we briefly recall the main features of the nuclear rainbow phenomenon. When the absorption is weak, the interference pattern of scattering waves attractively refracted by the target (serving as a nuclear liquid drop) turns out to be analogous to the interference of light rays refracted by a water droplet that causes the atmospheric rainbow. As schematically illustrated in Fig. 1, the trajectories incident at the sub-surface impact parameters are attractively refracted to the far side of the scattering center [4, 5], and the interference of the two subamplitudes of the farside scattering ($f_{F>}$ and $f_{F<}$) gives rise to a broad oscillation pattern of elastic scattering cross section, which is a microscopic replica of the Airy oscillation pattern of the refracted light rays of atmospheric rainbow [3, 6]. Thus, the key to identify nuclear rainbow is the Airy oscillation of the farside cross section in elastic nucleus-nucleus scattering, especially, the first Airy minimum A1 followed by a shoulder-like bump in elastic scattering cross section at medium and large angles [3, 6, 7].
@@
A more popular, alternative interpretation of the farside scattering is based on the decomposition of the elastic scattering amplitude into the nearside and farside components, using the method originally developed by Fuller [10] which is referred to hereafter as the nearside-farside (NF) decomposition. The Airy oscillation pattern of nuclear rainbow is formed by an interference of the outer ($f_{F>}$) and inner ($f_{F<}$) subamplitudes of the farside scattering. The $f_{F>}$ and $f_{F<}$ subamplitudes represent the contributions to the same scattering angle $\theta$ of the (farside) partial waves with angular momenta $\ell>\ell_{\rm R}$ and $\ell<\ell_{\rm R}$, respectively, where $\ell_{\rm R}$ is the angular momentum at the minimum of the deflection function, the rainbow point [11] that determines the corresponding rainbow angle $\Theta_{\rm R}$ in the observable angular range ($0<\Theta_{\rm R}<180^{\circ}$) [12]. The scattering to angles $\theta>\Theta_{\rm R}$, i.e., to the dark side of the rainbow, is classically forbidden. The location of $\Theta_{\rm R}$ is shifting from large scattering angles at low energies to medium angles of $\Theta_{\rm R}\sim 40^{\circ}-60^{\circ}$ at energies approaching Fermi domain of around 20 MeV/nucleon, where pronounced nuclear rainbow patterns have been observed [3]. It is complementary to note that $\Theta_{\rm R}\approx 138^{\circ}$ for the refracted light rays of atmospheric rainbow [4]. In elastic HI scattering, the $f_{F<}$ subamplitude of the farside scattering is often suppressed by the absorption, and the Airy oscillation pattern of nuclear rainbow disappears. Therefore, a weak absorption is a prerequisite for the formation and observation of nuclear rainbow.
@@
As mentioned above, the observation of the (refractive) nuclear rainbow pattern provides important database for studies of the nucleus-nucleus interaction at low and medium energies [2, 3]. Because nuclear rainbow is formed mainly by the farside trajectories refracted by a strongly attractive nucleus-nucleus OP, the nuclear rainbow pattern is absent in elastic electron-nucleus or nucleon-nucleus scattering where the scattering potential is not attractive enough to give rise to the farside scattering [6]. Thus, nuclear rainbow can appear only in the refractive $\alpha$-nucleus or nucleus-nucleus scattering which is governed by the deep attractive real OP [3].
The nuclear rainbow pattern has been shown mainly so far by using the Fuller decomposition method [10]. However, this method was formulated for a nonidentical dinuclear system only, and it remains unexplored how the projectile-target exchange symmetry of an identical system affects the nuclear rainbow pattern. For this purpose, we have generalized in the present work Fuller method for the elastic scattering of two identical (even-even) nuclei, taking into account exactly the projectile-target exchange symmetry. The generalized method is then used to study the nuclear rainbow pattern in elastic 12C+12C and 16O+16O scattering, which was proven to be strongly refractive [13, 14, 15, 16, 17, 3, 18].
@@ II NF decomposition of the elastic scattering amplitude of two identical spin-zero nuclei
We briefly recall how the nuclear rainbow pattern is revealed by decomposing the elastic scattering amplitude into the nearside $f^{(\rm N)}$ and farside $f^{(\rm F)}$ components using Fuller method [10]. By splitting the Legendre function $P_{\ell}(\cos\theta)$ into two waves scattered at $\theta$ but running in the opposite directions around the scattering center, the NF components of the unsymmetrized (nuclear) elastic scattering amplitude are obtained as
@@
Because of the short range of strong interaction, the NF decomposition of the nuclear scattering amplitude (3) can be done numerically, with the nuclear nearside and farside amplitudes merging naturally at $\theta=0$ and $\theta=\pi$. However, the long-range Coulomb interaction poses a technical difficulty for the NF decomposition of the Coulomb scattering amplitude based on the partial wave expansion (3). Moreover, the Coulomb scattering amplitude is singular at $\theta=0$, where the partial wave series diverges. By projecting argument of the Legendre function into the complex plane, Fuller managed to treat the singularity at $\theta=0$ and obtained the NF components $f^{\rm(N)}_{\rm R}(\theta)$ and $f^{\rm(F)}_{\rm R}(\theta)$ of the Rutherford scattering amplitude [10] as
@@
$S(\theta)=(1+i\eta)^{-1}F(1,1+i\eta;2+i\eta;\sin^{2}(\theta/2)),\mbox{and}\hskip 36.98866pt$
hypergeometric $\mbox{hypergeometric function}\ F(a,b;c;z)={}_{2}F_{1}(a,b;c;z)\hskip 36.98866pt$
$=\frac{\Gamma(c)}{\Gamma(a)\Gamma(b)}\sum_{n=0}^{\infty}\frac{\Gamma(a+n)\Gamma(b+n)}{\Gamma(c+n)}\frac{z^{n}}{z!}.\hskip 51.21504pt$ (6d)
@@ III Results for the elastic 12C+12C and 16O+16O scattering
Figure 4: (a) Elastic 12C+12C scattering data measured at $E_{\rm lab}=112$ MeV [23] in comparison with results of the OM calculation, taking exactly into account the projectile-target symmetrization (solid line). The nearside (dotted line) and farside (dashed line) cross sections were given by the NF decomposition (16)-(17) of the ES amplitude. Ak and A’k are the k-order Airy minimum of the direct farside cross section at $\theta$ and its symmetric partner of the exchange farside cross section at $\pi-\theta$, respectively.
@@
We have explored such a sensitivity to the real OP of the elastic 12C+12C data measured at $E_{\rm lab}=112$ MeV by slightly rescaling the strength of the best-fit real folded OP up and down by $\sim 7\%$, with the depth of the WS imaginary OP being adjusted in each case by $\chi^{2}$ fit to the measured data. The radial shapes of the complex 12C+12C OP are shown in Fig. 6, and the corresponding OM results are compared with the data in Fig. 7. It can be seen in panel (b) of Fig. 7 that the Airy pattern of the (direct) farside scattering cross section is very sensitive to the strength of the real OP. The rescaling of the real OP shown in Fig. 6 results in different locations of the Airy minima of the direct farside cross section, which strongly affect the interference of the direct and exchange farside amplitudes at angles around $90^{\circ}$, as shown in panel (a) of Fig. 7. Similar sensitivity of the elastic 12C+12C scattering data at $E_{\rm lab}=78$ MeV [23] to the real OP was used in a recent study of the 12C+12C fusion [28] to probe different treatments of the nuclear overlap density in the double-folding calculation of the real OP of the 12C+12C system at astrophysical energies.
@@ IV Elastic 16O+12C scattering and the core-core exchange effect
As variance with the total elastic scattering amplitude (9) of two identical nuclei, summation of the partial wave series (22) is done over both odd and even partial waves $\ell$, with the parity-dependent elastic $\alpha$ transfer $S_{\rm ET}$ matrix added to the elastic scattering $S_{\rm ES}$ matrix. The interference of $S_{\rm ES}$ and $S_{\rm ET}$ leads to the oscillation of elastic 16O+12C cross section observed at large angles (see Fig. 9). Schematically, the elastic $\alpha$ transfer shown in Fig. 8 can also be treated as the core-core exchange that naturally leads to that interference.
@@
Note that the ES and ET contributions to the total elastic 16O+12C cross section are not symmetrically equal as the direct and exchange scattering amplitudes found above for two identical nuclei. To explore this effect in more details, the NF decomposition of the elastic 16O+12C amplitude, given by the two-channel CRC calculation at $E_{\rm lab}=132$ MeV, has been done separately for the ES, ET, and total (ES+ET) elastic amplitude using Fuller’s method [10]. One can see in panels (b) of Fig. 10 that the ET cross section is not a symmetrically reflected pattern of the ES cross section shown in panels (a). While the ES cross section is dominated by the farside scattering over a wide angular range, the ET cross section is a typical NF interference pattern, which indicates a surface character of the ET between two 12C cores. This is natural because the contribution of $f_{\rm ET}(\pi-\theta)$ to $f_{\rm total}(\theta)$ at the most backward angles represents in fact the ET process occurring physically at the most forward angles.
As variance with the ES occurring at forward angles that includes both the Coulomb and nuclear scattering, the repulsive Coulomb interaction does not contribute to the ET process which occurs at backward angles. As a result, we found a significant non-refractive farside component of the ET cross section shown in panel (b) of Fig. 10. Although generated by the attractive nuclear interaction, such a farside cross section cannot be associated with nuclear rainbow due to a quick modulation of the ET farside cross section caused mainly by the diffraction of the $\alpha$-transfer wave.
@@
Like the total ES amplitude of two identical nuclei (16)-(17), the nearside component (23) of the total elastic 16O+12C amplitude is a superposition of the nearside and farside components of the ES and ET amplitudes determined at angles $\theta$ and $\pi-\theta$, respectively, and vice versa for the farside component (24) of the total elastic amplitude. At the most backward angles, the ET process become dominant and the nearside component of the total elastic amplitude is determined entirely by the farside component of the ET amplitude, and vice versa for the farside component of the total scattering amplitude, as shown in panels (b) and (c) of Fig. 10. Although the NF pattern is not symmetric as found for the elastic scatetring of two identical nuclei, the NF interchange between the total elastic and ET amplitudes is naturally caused by the ET between two identical 12C cores, i.e., the core-core exchange symmetry of the 16O+12C system.
@@ V Summary
A similar NF interchange of the total elastic and elastic $\alpha$ transfer amplitudes was found in the nonidentical 16O+12C system with the core-core symmetry, where the ET process becomes dominant at backwards angles. As variance with the pure elastic scattering, a significant non-refractive farside component of the ET cross section was found, which indicates the surface character of the $\alpha$ transfer between two 12C cores.

### v2

@@ Abstract
Extensive elastic scattering data measured at energies around 10 to 20 MeV/nucleon for some light identical systems, like 12C+12C and 16O+16O , were shown to exhibit the nuclear rainbow pattern of broad Airy oscillations of the elastic scattering cross section at medium and large angles. Because of the identity of the scattered projectile and recoiled target, the smooth rainbow pattern at angles around and beyond $\theta_{\rm c.m.}\approx 90^{\circ}$ is strongly deteriorated by the boson exchange in the 12C+12C and 16O+16O systems at low energies. The exchange symmetry of two identical nuclei implies the Mott interference of the direct and exchange scattering amplitudes, which destroys the nuclear rainbow pattern. The nuclear rainbow features in the elastic scattering of two identical nuclei have been discussed so far based on the nearside-farside (NF) decomposition of the scattering amplitude given by an optical model calculation neglecting the projectile-target exchange symmetry. Moreover, the NF decomposition method was developed in the 1970s by Fuller for nonidentical dinuclear systems only, and the details of how the exchange symmetry of an identical system affects the evolution of nuclear rainbow remain unexplored. For this purpose, the Fuller method is generalized in the present work for the elastic scattering of two identical (spin-zero) nuclei, with the projectile-target exchange symmetry taken explicitly into account. The results obtained for elastic 12C+12C and 16O+16O scattering at low energies show that the exchange symmetry results in a symmetric interchange of the nearside and farside scattering patterns at angles passing through $\theta_{\rm c.m.}=90^{\circ}$, which requires a more subtle interpretation of nuclear rainbow. We found further that a similar NF interchange also occurs in a nonidentical nucleus-nucleus system with the core-core symmetry at low energies, where the elastic cross section at backward angles is due mainly to the elastic transfer of cluster or nucleon between two identical cores. This interesting effect is illustrated in the elastic 16O+12C scattering at low energies where the elastic $\alpha$ transfer between two 12C cores has been proven to enhance the elastic cross section at backward angles.
@@ I Introduction
To motivate the present study, we briefly recall the main features of the nuclear rainbow phenomenon. When the absorption is weak, the interference pattern of scattering waves attractively refracted by the target (serving as a nuclear liquid drop) turns out to be analogous to the interference of light rays refracted by a water droplet that causes the atmospheric rainbow. As schematically illustrated in Fig. 1, the trajectories incident at the sub-surface impact parameters are attractively refracted to the far side of the scattering center [4, 5], and the interference of the two subamplitudes of the farside scattering ($f_{F>}$ and $f_{F<}$) gives rise to a broad oscillation pattern of elastic scattering cross section, which is a microscopic replica of the Airy oscillation pattern of the refracted light rays of atmospheric rainbow [3, 6]. Thus, the key to identify a nuclear rainbow is the Airy oscillation of the farside cross section in elastic nucleus-nucleus scattering, especially, the first Airy minimum A1 followed by a shoulder-like bump in elastic scattering cross section at medium and large angles [3, 6, 7].
@@
A more popular, alternative interpretation of the farside scattering is based on the decomposition of the elastic scattering amplitude into the nearside and farside components, using the method originally developed by Fuller [10] which is referred to hereafter as the nearside-farside (NF) decomposition. The Airy oscillation pattern of the nuclear rainbow is formed by an interference of the outer ($f_{F>}$) and inner ($f_{F<}$) subamplitudes of the farside scattering. The $f_{F>}$ and $f_{F<}$ subamplitudes represent the contributions to the same scattering angle $\theta$ of the (farside) partial waves with angular momenta $\ell>\ell_{\rm R}$ and $\ell<\ell_{\rm R}$, respectively, where $\ell_{\rm R}$ is the angular momentum at the minimum of the deflection function, the rainbow point [11] that determines the corresponding rainbow angle $\Theta_{\rm R}$ in the observable angular range ($0<\Theta_{\rm R}<180^{\circ}$) [12]. The scattering to angles $\theta>\Theta_{\rm R}$, i.e., to the dark side of the rainbow, is classically forbidden. The location of $\Theta_{\rm R}$ is shifting from large scattering angles at low energies to medium angles of $\Theta_{\rm R}\sim 40^{\circ}-60^{\circ}$ at energies approaching Fermi domain of around 20 MeV/nucleon, where pronounced nuclear rainbow patterns have been observed [3]. It is complementary to note that $\Theta_{\rm R}\approx 138^{\circ}$ for the refracted light rays of atmospheric rainbow [4]. In elastic HI scattering, the $f_{F<}$ subamplitude of the farside scattering is often suppressed by the absorption, and the Airy oscillation pattern of nuclear rainbow disappears. Therefore, a weak absorption is a prerequisite for the formation and observation of nuclear rainbow.
@@
As mentioned above, the observation of the (refractive) nuclear rainbow pattern provides an important database for studies of the nucleus-nucleus interaction at low and medium energies [2, 3]. Because a nuclear rainbow is formed mainly by the farside trajectories refracted by a strongly attractive nucleus-nucleus OP, the nuclear rainbow pattern is absent in elastic electron-nucleus or nucleon-nucleus scattering where the scattering potential is not attractive enough to give rise to the farside scattering [6]. Thus, the nuclear rainbow can appear only in the refractive $\alpha$-nucleus or nucleus-nucleus scattering which is governed by the deep attractive real OP [3].
The nuclear rainbow pattern has been shown mainly so far by using the Fuller decomposition method [10]. However, this method was formulated for a nonidentical dinuclear system only, and it remains unexplored how the projectile-target exchange symmetry of an identical system affects the nuclear rainbow pattern. For this purpose, we have generalized in the present work, the Fuller method for the elastic scattering of two identical (even-even) nuclei, taking into account exactly the projectile-target exchange symmetry. The generalized method is then used to study the nuclear rainbow pattern in elastic 12C+12C and 16O+16O scattering, which was proven to be strongly refractive [13, 14, 15, 16, 17, 3, 18].
@@ II NF decomposition of the elastic scattering amplitude of two identical spin-zero nuclei
We briefly recall how the nuclear rainbow pattern is revealed by decomposing the elastic scattering amplitude into the nearside $f^{(\rm N)}$ and farside $f^{(\rm F)}$ components using Fuller method [10]. By splitting the Legendre function $P_{\ell}(\cos\theta)$ into two waves scattered at $\theta$ but running in the opposite directions around the scattering center, the NF components of the unsymmetrized (nuclear) elastic scattering amplitude is obtained as
@@
Because of the short range of the strong interaction, the NF decomposition of the nuclear scattering amplitude (3) can be done numerically, with the nuclear nearside and farside amplitudes merging naturally at $\theta=0$ and $\theta=\pi$. However, the long-range Coulomb interaction poses a technical difficulty for the NF decomposition of the Coulomb scattering amplitude based on the partial wave expansion (3). Moreover, the Coulomb scattering amplitude is singular at $\theta=0$, where the partial wave series diverges. By projecting argument of the Legendre function into the complex plane, Fuller managed to treat the singularity at $\theta=0$ and obtained the NF components $f^{\rm(N)}_{\rm R}(\theta)$ and $f^{\rm(F)}_{\rm R}(\theta)$ of the Rutherford scattering amplitude [10] as
@@
$S(\theta)=(1+i\eta)^{-1}F(1,1+i\eta;2+i\eta;\sin^{2}(\theta/2)),\hskip 36.98866pt$ (6d)
and hypergeometric function
$F(a,b;c;z)={}_{2}F_{1}(a,b;c;z),\hskip 128.0374pt$
$=\frac{\Gamma(c)}{\Gamma(a)\Gamma(b)}\sum_{n=0}^{\infty}\frac{\Gamma(a+n)\Gamma(b+n)}{\Gamma(c+n)}\frac{z^{n}}{z!}.\hskip 35.56593pt$ (6e)
@@ III Results for the elastic 12C+12C and 16O+16O scattering
[…]
@@
We have explored such a sensitivity to the real OP of the elastic 12C+12C data measured at $E_{\rm lab}=112$ MeV by slightly rescaling the strength of the best-fit real folded OP up and down by $\approx 7\%$, with the depth of the WS imaginary OP being adjusted in each case by $\chi^{2}$ fit to the measured data. The radial shapes of the complex 12C+12C OP are shown in Fig. 6, and the corresponding OM results are compared with the data in Fig. 7. It can be seen in panel (b) of Fig. 7 that the Airy pattern of the (direct) farside scattering cross section is very sensitive to the strength of the real OP. The rescaling of the real OP shown in Fig. 6 results in different locations of the Airy minima of the direct farside cross section, which strongly affect the interference of the direct and exchange farside amplitudes at angles around $90^{\circ}$, as shown in panel (a) of Fig. 7. Similar sensitivity of the elastic 12C+12C scattering data at $E_{\rm lab}=78$ MeV [23] to the real OP was used in a recent study of the 12C+12C fusion [28] to probe different treatments of the nuclear overlap density in the double-folding calculation of the real OP of the 12C+12C system at astrophysical energies.
@@ IV Elastic 16O+12C scattering and the core-core exchange effect
At variance with the total elastic scattering amplitude (9) of two identical nuclei, summation of the partial wave series (22) is done over both odd and even partial waves $\ell$, with the parity-dependent elastic $\alpha$ transfer $S_{\rm ET}$ matrix added to the elastic scattering $S_{\rm ES}$ matrix. The interference of $S_{\rm ES}$ and $S_{\rm ET}$ leads to the oscillation of elastic 16O+12C cross section observed at large angles (see Fig. 9). Schematically, the elastic $\alpha$ transfer shown in Fig. 8 can also be treated as the core-core exchange that naturally leads to that interference.
@@
Note that the ES and ET contributions to the total elastic 16O+12C cross section are not symmetrically equal as the direct and exchange scattering amplitudes found above for two identical nuclei. To explore this effect in more details, the NF decomposition of the elastic 16O+12C amplitude, given by the two-channel CRC calculation at $E_{\rm lab}=132$ MeV, has been done separately for the ES, ET, and total (ES+ET) elastic amplitude using Fuller’s method [10]. One can see in panel (b) of Fig. 10 that the ET cross section is not a symmetrically reflected pattern of the ES cross section shown in panels (a). While the ES cross section is dominated by the farside scattering over a wide angular range, the ET cross section is a typical NF interference pattern, which indicates a surface character of the ET between two 12C cores. This is natural because the contribution of $f_{\rm ET}(\pi-\theta)$ to $f_{\rm total}(\theta)$ at the most backward angles represents in fact the ET process occurring physically at the most forward angles.
At variance with the ES occurring at forward angles that includes both the Coulomb and nuclear scattering, the repulsive Coulomb interaction does not contribute to the ET process which occurs at backward angles. As a result, we found a significant non-refractive farside component of the ET cross section shown in panel (b) of Fig. 10. Although generated by the attractive nuclear interaction, such a farside cross section cannot be associated with nuclear rainbow due to a quick modulation of the ET farside cross section caused mainly by the diffraction of the $\alpha$-transfer wave.
@@
Like the total ES amplitude of two identical nuclei (16)-(17), the nearside component (23) of the total elastic 16O+12C amplitude is a superposition of the nearside and farside components of the ES and ET amplitudes determined at angles $\theta$ and $\pi-\theta$, respectively, and vice versa for the farside component (24) of the total elastic amplitude. At the most backward angles, the ET process become dominant and the nearside component of the total elastic amplitude is determined entirely by the farside component of the ET amplitude, and vice versa for the farside component of the total scattering amplitude, as shown in panels (b) and (c) of Fig. 10. Although the NF pattern is not symmetric as found for the elastic scattetring of two identical nuclei, the NF interchange between the total elastic and ET amplitudes is naturally caused by the ET between two identical 12C cores, i.e., the core-core exchange symmetry of the 16O+12C system.
@@ V Summary
A similar NF interchange of the total elastic and elastic $\alpha$ transfer amplitudes was found in the nonidentical 16O+12C system with the core-core symmetry, where the ET process becomes dominant at backwards angles. At variance with the pure elastic scattering, a significant non-refractive farside component of the ET cross section was found, which indicates the surface character of the $\alpha$ transfer between two 12C cores.